# Recording for Eyes, Not Echoing to Ears: Contextualized Spoken-to-Written Conversion of ASR Transcripts

### v2

@@ Introduction
As can be seen in Figure 1, since ASR transcripts aim to provide verbatim transcriptions of oral communications, they often exhibit various spoken language phenomena and informal styles, such as filler words, repetitions, repairs, and fragments, and include ASR errors and ungrammatical text. These characteristics lead to poor readability. To enhance the readability of ASR transcripts, we propose the Contextualized Spoken-to-Written conversion (CoS2W) task, which aims to correct ASR errors and grammatical errors and transfer the informal style to the written and formal style while preserving the content, that is, the CoS2W task is designed to convert verbatim transcripts (“echoing to ears”) into readable documents (“recording for eyes”). Note that different domains vary in their requirements for adjustments on ASR transcripts. Some domains only allow removal of filler words while some domains (e.g., speech analysis for diagnosis) require verbatim transcripts. Nonetheless, the readable documents of CoS2W task are particularly helpful for many domains such as podcast, education, and meetings (business, project, etc.), where efficiency in information delivery and knowledge acquisition is essential.
Moreover, for a wide variety of downstream tasks on ASR transcripts, such as machine translation, summarization, question-answering (Gupta et al. 2021), and sentiment classification (Zhang et al. 2023b), we find that even competitive models (including LLMs) perform substantially better on written text than on ASR transcripts, and the performance on the downstream tasks could benefit substantially from applying CoS2W to ASR transcripts. For example, for machine translation of ASR transcripts, we employ the competitive CSANMT model  [^1] https://huggingface.co/modelscope-unofficial/damo-csanmt-zh-en-large on the development set of the speech translation BSTC dataset (Zhang et al. 2021b) and find that compared to directly translating ASR transcripts, applying CoS2W (with GPT-4) on ASR transcripts then translating improves the BLEURT (Sellam, Das, and Parikh 2020) score from 57.07 to 61.9.
@@
Grammar Error Correction task entails correcting textual grammatical mistakes. Similar to ASR error correction, LLM performance in GEC isn’t fully satisfactory (Zhang et al. 2023a; Fang et al. 2023). GEC datasets usually follow minimum change principle, clashing with LLMs’ relatively free generation paradigm. Disregarding objective metrics, human evaluations confirm LLMs’ GEC effectiveness (Li et al. 2023), which suggests current metrics don’t evaluate LLMs fairly and reliably. Therefore, Li et al. (2024) aim to refine GEC evaluation methods, using LLMs to categorize edits and calculate metrics like $F_{0.5}$ to evaluate models.
@@
To conduct experiments and evaluations of the CoS2W task, we construct the Spoken2Written of ASR Transcripts Benchmark (SWAB) dataset, including multiple scenarios (i.e., podcasts, meetings, lectures) in both Chinese and English. There are 60 transcripts with auxiliary information, with each subcategory comprising 10 documents. It is important to note that the SWAB is document-level with complete annotations for entire transcripts, providing an annotated target as a reference for each paragraph. Furthermore, we provide links to the original audio or video along with timestamps to support multi-modal research.
@@
Lectures are sourced from YiXi speeches for Chinese and TED talks for English on YouTube, covering a variety of topics. Both Chinese and English lectures are provided with meta-information of the speeches as auxiliary information, which includes the name and biography of the speaker, as well as the title, category, and description of the video.
@@
All collected data are first transcribed using a competitive ASR system (Gao et al. 2022)  [^3] https://github.com/modelscope/FunASR, then the ASR 1-bests are processed with punctuation insertion, paragraph segmentation (Zhang et al. 2021a), and speaker diarization (Zheng et al. 2023). This ASR structure is widely utilized and operates at peak performance while preserving the overall generality of the dataset. We utilize GPT-4 (gpt-4-0125-preview)  [^4] https://platform.openai.com/docs/models/ to obtain the initial target of CoS2W for the SWAB dataset. The input is chunk-level to leverage local context and to reduce token consumption. The auxiliary information is also provided to enhance the ASR error correction performance. The prompt for this procedure is in the appendix.
@@
LLaMA3-8B None 8.94 20.73 40.53 13.18 6.35 7.33
QWen-14B None 3.41 15.96 30.14 11.46 3.72 6.04
QWen-Max None 9.47 28.92 50.58 22.35 6.00 7.55
@@
To assess different LLMs’ abilities in solving this complex task, we choose 4 typical LLMs across both open-source and closed-source models. The open-source LLMs selected are QWen-14B  [^5] https://www.modelscope.cn/models/qwen/Qwen-14B-Chat and LLaMA3-8B  [^6] https://github.com/meta-llama/llama, chosen for their suitable computational demands for future SFT-related work. For closed-source LLMs, we pick GPT4 (gpt-4-0125-preview)  [^7] https://platform.openai.com/docs/models/ and QWen-Max (qwen-max-0107)  [^8] https://help.aliyun.com/zh/dashscope/developer-reference/, to obtain the state-of-the-art performance of LLMs on this task.
@@
As previously mentioned, auxiliary information can enhance ASR error correction by addressing challenging pronunciations. We utilize auxiliary information by one-step and two-stage methods. The one-step method integrates full or retrieved auxiliary information as a reference part of the context in the prompt. The retrieved part is top-K relevant sentences of auxiliary information for current content based on BM25. For the two-stage approach, we first use LLMs to extract the summary or keywords from the auxiliary information, with the prompts detailed in the appendix. We require the summary to retain all key terms as much as possible, especially challenging words for ASR. The keywords should be categorized by their entity types. Next, we use either the summary or keywords as references. And we require LLMs to avoid over-correction and unfaithful information injection in the prompt shown in the appendix.
@@
In addition, we utilize an LLM (i.e., GPT4) as the evaluator. Using the prompts shown in the appendix, we request the LLM to evaluate and score the model-generated target across faithfulness (S-Faithful) and formality (S-Formal). Scores range from 1 (worst) to 10 (best) for each criterion. For chunk-level results, we require the outputs to align each paragraph by the paragraph index, so we evaluate them at the paragraph-level granularity. To fairly compare paragraph-level and chunk-level results, we evaluate them both at paragraph and chunk levels. To enhance effectiveness, we sample a maximum of 100 paragraphs per document for evaluation. Furthermore, we sample over 200 paragraphs, whose results of different LLMs are ranked by human evaluators.
@@ Results
GPT4 (paragraph) 340.50 63.32 0.00
   w/ local context 654.47 67.11 0.00
GPT4 (chunk) 1515.00 62.15 1.60
   w/ auxiliary 2006.86 62.73 1.28
GPT4 (document) 8140.42 10.64 76.10
QWen-14B (paragraph) 340.50 59.87 0.00
   w/ local context 654.47 55.64 0.00
@@
Formality Score: As shown in Figure 3, the Spearman’s correlation coefficient between the LLM evaluation and human evaluation is 0.61. Although there is a gap compared to human performance, there is also a strong correlation, which validates that the formality scores are reliable. Moreover, we compare the commonly used evaluation method. Using a RoBERTa-based classifier  [^9] https://huggingface.co/s-nlp/roberta-base-formality-ranker, we classify English paragraphs as formal or informal and assign the formality probability as the formality score. The LLM evaluator outperforms the classifier in the correlation coefficient, confirming the reliability of LLM evaluation.
As Figure 3 illustrates, GPT-4 Evaluator correlates more strongly with human evaluations than QWen-Max Evaluator, confirming its reliability. In addition, we evaluate GPT-4 results three times. The average standard deviations of faithfulness scores (0.19) and formality scores (0.11) across all samples confirms the robustness. Despite LLM updates rapidly, stable evaluation results can still be achieved with the same model version at a low temperature.
@@
As indicated in Table 6, LLMs generally excel more in Chinese than English at the chunk level. There are more blank paragraphs in English (2.68% for GPT4) than in Chinese (0.83%). This may be because LLMs prioritize fluency in English over Chinese, often merging paragraphs for better fluency at the expense of faithfulness.
@@ Appendix F ASR Performance
Appendix G Reproducibility Checklist
This paper:
  • Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes)
  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes)
  • Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper (yes)
Does this paper make theoretical contributions? (yes)
  • All assumptions and restrictions are stated clearly and formally. (yes)
  • All novel claims are stated formally (e.g., in theorem statements). (yes)
  • Proofs of all novel claims are included. (yes)
  • Proof sketches or intuitions are given for complex and/or novel results. (yes)
  • Appropriate citations to theoretical tools used are given. (yes)
  • All theoretical claims are demonstrated empirically to hold. (yes)
  • All experimental code used to eliminate or disprove claims is included. (no: The related code will be made publicly available upon publication of the paper.)
Does this paper rely on one or more datasets? (yes)
  • A motivation is given for why the experiments are conducted on the selected datasets (yes)
  • All novel datasets introduced in this paper are included in a data appendix. (yes)
  • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes)
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations. (yes)
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available. (yes)
  • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing. (yes)
Does this paper include computational experiments? (yes)
  • Any code required for pre-processing data is included in the appendix. (partial: We provide the prompt for dataset construction, and the related code will be made publicly available upon publication of the paper.)
  • All source code required for conducting and analyzing the experiments is included in a code appendix. (partial: We provide the prompts for the experiments and evaluations in the appendix, and the related code will be made publicly available upon publication of the paper.)
  • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes. (yes)
  • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (yes)
  • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results. (partial: We present the configurations for LLMs, while the generated results still demonstrate a degree of randomness.)
  • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks. (yes)
  • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics. (yes)
  • This paper states the number of algorithm runs used to compute each reported result. (yes)
  • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information. (yes)
  • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank). (yes)
  • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (partial)
  • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (yes)

### v4

@@ Introduction
As can be seen in Figure 1, since ASR transcripts aim to provide verbatim transcriptions of oral communications, they often exhibit various spoken language phenomena and informal styles, such as filler words, repetitions, repairs, and fragments, and include ASR errors and ungrammatical text. These characteristics lead to poor readability. To enhance the readability of ASR transcripts, we propose the Contextualized Spoken-to-Written conversion (CoS2W) task, which aims to correct ASR errors and grammatical errors and transfer the informal style to the written and formal style while preserving the content, that is, the CoS2W task is designed to convert verbatim transcripts (“echoing to ears”) into readable documents (“recording for eyes”). Note that different domains vary in their requirements for adjustments on ASR transcripts. Some domains only allow removal of filler words while some domains (e.g., speech analysis for diagnosis) require verbatim transcripts. Nonetheless, the readable documents of CoS2W task are particularly helpful for many domains such as podcasts, education, and meetings (business, project, etc.), where efficiency in information delivery and knowledge acquisition is essential.
Moreover, for a wide variety of downstream tasks on ASR transcripts, such as machine translation, summarization, question-answering (Gupta et al. 2021), and sentiment classification (Zhang et al. 2023), we find that even competitive models (including LLMs) perform substantially better on written text than on ASR transcripts, and the performance on the downstream tasks could benefit substantially from applying CoS2W to ASR transcripts. For example, for machine translation of ASR transcripts, we employ the competitive CSANMT model [^1] https://www.huggingface.co/modelscope-unofficial/damo-csanmt-zh-en-large on the development set of the speech translation BSTC dataset (Zhang et al. 2021b) and find that compared to directly translating ASR transcripts, applying CoS2W (with GPT-4) on ASR transcripts then translating improves the BLEURT (Sellam, Das, and Parikh 2020) score from 57.07 to 61.9.
@@
Grammar Error Correction task entails correcting textual grammatical mistakes. Similar to ASR error correction, LLM performance in GEC isn’t fully satisfactory (Zhang et al. 2023; Fang et al. 2023). GEC datasets usually follow minimum change principle, clashing with LLMs’ relatively free generation paradigm. Disregarding objective metrics, human evaluations confirm LLMs’ GEC effectiveness (Li et al. 2023), which suggests current metrics don’t evaluate LLMs fairly and reliably. Therefore, Li et al. (2024) aim to refine GEC evaluation methods, using LLMs to categorize edits and calculate metrics like $F_{0.5}$ to evaluate models.
@@
To conduct experiments and evaluations of the CoS2W task, we construct the Spoken2Written of ASR Transcripts Benchmark (SWAB) dataset [^2] https://www.github.com/alibaba-damo-academy/SpokenNLP/tree/main/swab, including multiple scenarios (i.e., podcasts, meetings, lectures) in both Chinese and English. There are 60 transcripts with auxiliary information, with each subcategory comprising 10 documents. It is important to note that the SWAB is document-level with complete annotations for entire transcripts, providing an annotated target as a reference for each paragraph. Furthermore, we provide links to the original audio or video along with timestamps to support multi-modal research.
@@
Lectures are sourced from many Chinese and English individual speeches on YouTube, covering a variety of topics. Both Chinese and English lectures are provided with meta-information of the speeches as auxiliary information, which includes the name and biography of the speaker, as well as the title, category, and description of the video.
@@
All collected data are first transcribed using a competitive ASR system (Gao et al. 2022) [^4] https://www.github.com/modelscope/FunASR, then the ASR 1-bests are processed with punctuation insertion, paragraph segmentation (Zhang et al. 2021a), and speaker diarization (Zheng et al. 2023). This ASR structure is widely utilized and operates at peak performance while preserving the overall generality of the dataset. We utilize GPT-4 (gpt-4-0125-preview) [^5] https://platform.openai.com/docs/models to obtain the initial target of CoS2W for the SWAB dataset. The input is chunk-level to leverage local context and to reduce token consumption. The auxiliary information is also provided to enhance the ASR error correction performance. The prompt for this procedure is in the appendix.
@@
LLaMA3-8B None 08.94 20.73 40.53 13.18 6.35 7.33
QWen-14B None 03.41 15.96 30.14 11.46 3.72 6.04
QWen-Max None 09.47 28.92 50.58 22.35 6.00 7.55
@@
To assess different LLMs’ abilities in solving this complex task, we choose 4 typical LLMs across both open-source and closed-source models. The open-source LLMs selected are QWen-14B [^6] https://www.modelscope.cn/models/qwen/Qwen-14B-Chat and LLaMA3-8B [^7] https://www.github.com/meta-llama/llama, chosen for their suitable computational demands for future SFT-related work. For closed-source LLMs, we pick GPT-4 (gpt-4-0125-preview) [^8] https://platform.openai.com/docs/models and QWen-Max (qwen-max-0107) [^9] https://www.alibabacloud.com/help/en/model-studio/getting-started/models, to obtain the state-of-the-art performance of LLMs on this task.
@@
We utilize auxiliary information to enhance ASR error correction by one-step and two-stage methods. The one-step method integrates full or retrieved auxiliary information as a reference part of the context in the prompt. The retrieved part is top-K relevant sentences of auxiliary information for current content based on BM25. For the two-stage approach, we first use LLMs to extract the summary or keywords from the auxiliary information, with the prompts detailed in the appendix. We require the summary to retain all key terms as much as possible, especially challenging words for ASR. The keywords should be categorized by their entity types. Next, we use either the summary or keywords as references.
@@
In addition, we utilize an LLM (i.e., GPT-4) as the evaluator with the prompts shown in the appendix. We request the LLM to evaluate and score the model-generated target across faithfulness (S-Faithful) and formality (S-Formal). Scores range from 1 (worst) to 10 (best) for each criterion. S-Faithful reflects content preservation to retain original meaning, and S-Formal reflects the degree of formality in style. For chunk-level results, we require the outputs to align each paragraph by the paragraph index, so we evaluate them at the paragraph-level granularity. To fairly compare paragraph-level and chunk-level results, we evaluate them both at paragraph and chunk levels. To enhance effectiveness, we sample a maximum of 100 paragraphs per document for evaluation. Furthermore, we sample over 200 paragraphs, whose results of different LLMs are ranked by human evaluators.
@@ Results
GPT-4 (paragraph) 0340.50 63.32 00.00
   w/ local context 0654.47 67.11 00.00
GPT-4 (chunk) 1515.00 62.15 01.60
   w/ auxiliary 2006.86 62.73 01.28
GPT-4 (document) 8140.42 10.64 76.10
QWen-14B (paragraph) 0340.50 59.87 00.00
   w/ local context 0654.47 55.64 00.00
@@
Formality Score: As shown in Figure 3, the Spearman’s correlation coefficient between the LLM evaluation and human evaluation is 0.61. Although there is a gap compared to human performance, there is also a strong correlation, which validates that the formality scores are reliable. Moreover, we compare the commonly used evaluation method. Using a RoBERTa-based classifier [^10] https://www.huggingface.co/s-nlp/roberta-base-formality-ranker, we classify English paragraphs as formal or informal and assign the formality probability as the formality score. The LLM evaluator outperforms the classifier in the correlation coefficient, confirming the reliability of LLM evaluation.
As Figure 3 illustrates, GPT-4 Evaluator correlates more strongly with human evaluations than QWen-Max Evaluator, confirming its reliability. In addition, we evaluate GPT-4 results three times. The average standard deviations of faithfulness scores (0.19) and formality scores (0.11) across all samples confirm the robustness. Despite LLM updates rapidly, stable evaluation results can still be achieved with the same model version at a low temperature.
@@
As indicated in Table 6, LLMs generally excel more in Chinese than English at the chunk level. There are more blank paragraphs in English (2.68% for GPT-4) than in Chinese (0.83%). This may be because LLMs prioritize fluency in English over Chinese, often merging paragraphs for better fluency at the expense of faithfulness.
@@ Conclusion
Ethics Statement
The SWAB dataset used in this research is strictly for academic and non-commercial purposes. We implemented several measures to ensure compliance with ethical standards, as follows.
  • Data Transparency and Anonymization. The Chinese meetings in the SWAB are sourced from the training set of the publicly available AliMeeting dataset, and the English meetings come from the publicly available AMI meeting dataset. For our collected podcasts and lectures, we only provide ASR transcripts after rigorous text anonymization processes and our annotations, to ensure transparency regarding the data sources and their usage while maintaining anonymity.
  • Data Access Compliance. To further ensure the ethical use of the dataset, we require researchers to contact us via email to confirm their compliance with ethical guidelines and the conditions outlined in our data usage declaration, before granting them access to the dataset. This procedure includes ensuring that they are aware of and adhere to the Personal Information Protection Law (PIPL) and any relevant legal frameworks regarding personal data usage.
  • Authorization. Any personal data should be used only with express authorization, ensuring lawful and fair processing in accordance with applicable laws.
Acknowledgments
We would like to express gratitude to the anonymous reviewers for their valuable feedback.